OpenStudy (anonymous):

what is the correct solution for this compound inequality. 19 < k + 9 ≤ 100

pooja195 (pooja195):

19 < k + 9 ≤ 100 We want to isolate the k so just subtract 9 from everything 19-9=? 9-9=? 100-9=?

pooja195 (pooja195):

Thats all it is.